OpenStudy (anonymous):

Find the specified term of each geometric sequence. a1=7, r=2;a4 is the answer 13? or 56?

OpenStudy (tkhunny):

a1 = 7 a2 = 7*2 a3 = 7*2^2 a4 = 7*2^3